OpenStudy (anonymous):

the difference between the reciprocal of two consecutive positif even number is 1/12. Find the two number.

OpenStudy (anonymous):

4,6

OpenStudy (anonymous):

1/x - 1/(x+2) =1/12 or x(x+2)=24 or x^2 +2x -24=0 or x^2 +6x-4x-24=0 or x(x+6) -4 (x+6)=0 or (x+6)(x-4)=0 hence x=4 or x=-6 thus the two +ve numbers are 4 and +6